# On the correlation between Architectural Smells and Static Analysis Warnings

### v1

@@ 3.2 Study Context
We selected projects from the Qualitas Corpus (QCD) collection of software systems [47], using the compiled version of the QCD [22]. Figure 1 shows the word cloud composed by the names of the 112 java projects in the dataset crafted by Tempero et al. [22]. Table 2 details metrics on the QCD.
Figure 1: Qualitas Corpus WordCloud [22].
@@
Lines of Code (LOC) $>$ 18 million
@@ 3.3 Study setup and data collection
    We analyzed the QCD with four SATs and averaged the results for project packages;
@@
    To get our study results, we analyzed the finalized dataset, comprising all SAW detected by the three SATs and the AS detected by the Arcan tool grouped by software package.
We analyzed the entire data set (112 Java projects) from the Qualitas Corpus [47]. However, we obtained results for only 103 projects because the computational time for the remaining ones would have been excessively long.
@@ 3.4 Data Analysis
We considered the classes affected by at least an AS and one static analysis warning at the class or package levels. One or more static analysis warnings could infect the same set of classes for each AS. This is because an AS may involve more than one class, while static analysis warnings involve only one class. In the case of projects not infected by static analysis warnings or AS, we did not consider them for the analysis.
To answer RQ1, we collected the QCD projects and ran the SAT and ARCAN. We merged the outputs of the tool. In our context, the dependent variable is the AS, and the independent variables are the SAW. To compute the correlation coefficients, we need to assess the distributions of our variable. Therefore we conjectured one hypothesis ($H_{\mathcal{N}}$) as follows:
  • $H_{1\mathcal{N}}$: The data do not follow a normal distribution.
@@
  • $H_{0\mathcal{N}}$: The data follow a normal distribution.
We tested $H_{\mathcal{N}}$ with the Anderson-Darling (AD) test [48]. The AD test assesses whether data samples derive from a specific probability distribution, such as the normal distribution. AD measures the difference between the sample data and the expected values from the tested distribution. More specifically, it evaluates differences in the cumulative distribution function (CDF) between the observed data and the hypothesized distribution [48].
The AD test and the Shapiro-Wilk (SW) test [49] are both statistical tests used to assess the normality of data.
The SW test focuses on the correlation between the observed data and the expected values under a normal distribution, emphasizing the smallest and largest values in the dataset. Therefore, according to Mishra et al. [50], SW is more appropriate when targeting small datasets ($\leq 50$ samples). On the other hand, the AD test considers a wider range of values, including those in the middle of the distribution, providing a more sensitive evaluation of normality, especially for larger sample sizes than our own. Therefore, we preferred AD over SW [51]. AD is also considered one of the most powerful statistical tools for detecting most deviations from normality [51, 52].
Test results allowed us to reject the null hypothesis for each test case, $H_{0\mathcal{N}}$, asserting that the data was not normally distributed. Therefore, we chose Spearman’s $\rho$ [41] to compute the correlation in place of Pearsons, which requires a normal distribution [53]. The non-parametric Spearman’s rank correlation coefficient, denoted by $\rho$, measures statistical dependence between two variables without necessarily assuming a linear relationship or even normality of those variables. Spearman’s Rho measures the strength and direction of the monotonic relationship between the ranks of paired data points. Table 3 presents $\rho$ values interpretation according to Dancey and Reidy [54].
@@
To answer RQ2, we computed Spearman’s $\rho$ for each SAW-AS combination. We used the interquartile ratio (IQR) method [55] to select the top correlated SAW for each SAT and AS, i.e., we computed the quartiles. We selected only the SAW with their Spearman’s $\rho$ value higher or equal to the third quartile (Q3), thus selecting the top 25%. We computed the $\mathcal{P}$ as follows: Let $\Sigma_{\text{SAT}_{i}}$ represent the set of Static Analysis Warnings specific to a particular Static Analysis Tool $\text{SAT}_{i}$. Then, $\sigma\in\Sigma_{\text{SAT}_{i}}$ signifies that $\sigma$ belongs to the set of SAWs specific to $\text{SAT}_{i}$; thus $\sigma_{j}$ is a specific SAW and $\alpha\in AS$ represents a specific AS.. Therefore, we define the relative frequency $\mathcal{P}(\sigma_{j},\alpha)$ as:
@@
  • The numerator represents the number of times the specific SAW, $\sigma_{j}\in\Sigma_{\text{SAT}_{i}}$, appears within the given AS, $\alpha$.
  • The denominator represents the total count of all SAWs belonging to a specific SAT, $\sigma$, in the given AS, $\alpha$.
More specifically, we computed $\mathcal{P}$ per each possible combination of AS. Therefore our data computed a $\mathcal{P}$-value for the 3 ASs, the healthy instances, and for combinations of ASs. We tested $H_{2}$, and $H_{3}$ via the WPT [56], which is a non-parametric statistical test that compares two related samples or paired data. WPT uses the absolute difference between the two observations to classify and then compare the sum of the positive and negative differences. The test statistic is the lowest of both. We selected WPT to test $H_{2}$ and $H_{3}$ because the data was not normally distributed; hence, we used it instead of the paired t-test, which assumes a normal data distribution.
Moreover, due to the high number of statistical tests performed, we needed to apply p-value correction. Correcting p-values is essential in multiple hypothesis testing to control the increased risk of Type I errors. When multiple tests are conducted simultaneously, the probability of incorrectly rejecting at least one true null hypothesis increases. A classical p-value correction technique is the Bonferroni correction. However, Bonferroni is known to offer a conservative correction [57]. Therefore, we decided to use the Benjamini & Hochberg (BH) p-value correction [58], which uses the false discovery rate (FDR) when performing multiple hypothesis tests as follows [59]:
@@
    Rank the p-values: Sort the p-values obtained from multiple tests in ascending order, denoted as $p_{(1)}\leq p_{(2)}\ldots\leq p_{(m)}$ for their respectively defined hypotheses $H_{(1)},H_{(2)},\ldots,H_{(m)}$.
@@
    where $m$ is the total number of tests, $q_{(i)}$ is the corrected probability for the hypothesis at moment $i$ and $p_{(j)}$ given the order rank $j\in\{1,\ldots,m\}$ of p-values, which stands as the probability at the rank position $j$.
@@
    Determine correction index: The detailed procedure identified the largest index $k$ for which BH holds and rejects null hypotheses as follows:
@@
    which is automatically computed with existing analytical software.
@@
    to the desired critical value, $\alpha$ (e.g., 0.05), to determine which hypotheses to reject.
Since we performed the BH correction on the p-values of the defined hypotheses, we maintained the standard value of the critical value $\alpha$ on 0.05, hence gaining power on controlling for error Type I and Type II [59].
To answer RQ3, we leverage our definition of $\mathcal{P}$, the default severity associated with the specific SAW, and the severity associated with the specific AS to build three different rankers. The ranking idea stemmed from previous research by Çarka et al. [31], in which PofB and PopT metrics were normalized and used to evaluate classifier prediction concerning inspection effort (see Section 2.2).
@@
  • SAW severity: we used SonarQube severity associated with each SAW.
  • $\mathcal{P}$: we used $\mathcal{P}$ and created separate rankings one for each AS combination.
  • AS severity: we used the known AS severity to order the data.
We note that the AS severity-based order is a special case; by design, that ranker is the optimal one because it uses the AS severity. RQ3 aims to use the other two rankers to prioritize those SAWs that are more prone to induce ASs. Therefore, the optima ranker is needed to compare the results of the other two. According to the interpretation of PofB and PopT, we selected the first xof the data from the three rankings. Therefore, for each selected percentage of data, we computed how many Critical, High, and Medium AS-prone SAWs we would capture.
3.5 Replicability
We provide a replication package containing the raw data, including the instructions for the SAT execution, the list of projects infected by SAW and AS, and the full statistical test results for both the normality hypothesis and for the RQ’s hypotheses  [^5] https://zenodo.org/doi/10.5281/zenodo.11366846.
@@ 4.1 Static Analysis Warnings and Architectural Smell Correlation (RQ1)
We tested $H_{1}$ with Spearman’s $\rho$. We excluded all cross-correlations among SAWs and focused on the correlations between SAW and AS. Figure 3 presents the distribution of Spearman’s $\rho$ averaged for SAW by AS and SAT. We can reject the null hypothesis for each of the 663 cases. Hence, we can affirm that there is a statistically significant correlation between the presence of static analyzer warnings and the presence of architectural smells. More specifically, SAWs were weak to moderately correlated to AS for each AS across SAT.
Figure 3: Spearman’s $\rho$ average distribution for SAW by AS and SAT (RQ1)
@@ 4.2 Static Analysis Warnings inducing Architectural Smell (RQ2)
We tested $H_{2}$ with WPT. We rejected the null hypothesis for 243,258 out of 275,380 cases, i.e., 88%. Therefore, we can, on average, affirm that there is a statistically significant difference in the co-occurrence of pairs of SAW and AS. More specifically, Figure 4 presents the distribution of $H_{2}$ test results. It is worth noticing that in the case of UD, there was almost an equal probability of rejecting or accepting the null hypothesis across SAWs. Moreover, we note that there is an equal probability of rejecting the null hypothesis in the case of Healthy cases as well as those affected by all of the three ASs.
Furthermore, we tested $H_{3}$ with WPT. We can reject the null hypothesis in 30 out of 42 cases, i.e., 72% of the cases. Table 4 shows the test cases for which we could not reject the null hypothesis. Therefore, we can affirm that there is a statistically significant difference in the co-occurrence of SAT-specific SAWs and ASs presence.
@@
More specifically, Table 5 presents the top five SAW co-occurring with specific AS. We selected the top five SAWs that co-occur the most for each AS, including the “healthy carriers”. We note that the most co-occurring SAWs belong to FindBugs. Similarly, FindBugs also has the most “healthy carriers” co-occurring SAWs. Moreover, where three ASs affect a single package, the FindBugs SAW co-occurs each time. Futhermore, the least co-occurring AS is UD.
Table 5: The top five SAW co-occurring with specific AS
SAT SAW AS % co-occurring
@@
FindBugs NM_METHOD_NAMING_CONVENTION Healthy 89,865
@@ 4.3 Architectural Smell Prioritization (RQ3)
Figure 5 presents the distribution of $\mathcal{P}$ averaged among SAW across SAT per AS. Therefore, on average, Figure 5 presents the percentage of occurrences of SAWs related to the occurrence of one or more ASs or their absence (i.e., healthy). On average, SAWs belonging to SonarQube are more prone to raise warnings on healthy packages and classes. Most SAWs co-occur with CD AS and have a similar presence among the SATs. Moreover, second to CD, SAWs co-occurred with all three ASs. In this context, FindBugs’ SAWs exhibit an average higher co-occurrence.
Figure 5: Distribution of $\mathcal{P}$ averaged among SAW across SAT per AS (RQ2)
Figure 6 presents the SAW prioritization based on the three rankings, i.e., SAW Severity-based, $\mathcal{P}$-based, and optimal. We computed how many medium, high, or critical AS could be discovered by inspecting the first xof data ranked by one of the three methods. According to its design, the optimal ranking prioritized the SAW co-occurring with high and critical AS. The SAW severity-based and UD+CD rankers demonstrate the most balanced performance, while the ranking based on SAW co-occurring with CD favors moderately severe AS.
Finally, we tested $H_{4}$ with the WPT. We could reject the null hypothesis in all cases for high and critical AS severities, in which the optimal ranker was compared with the SAW-severity-based and the $\mathcal{P}$ based rankers. Conversely, we could reject the hypothesis when comparing the SAW-severity-based and the $\mathcal{P}$ based rankers but only the specific case when comparing the UD + HL + CD proneness based on $\mathcal{P}$. Futhermore, when the difference was statically significant, the gains were in favor of the optimal rankers, as expected.
We included the full table and the graphical representation of the gains among the approaches in the replication package for space constraint.
Moreover, we can not reject the null hypothesis for the medium risk for each ranker. Therefore, the optimal ranker behaves similarly to the other two rankers.
@@
Nevertheless, based on the results, we can affirm that SAW prioritization based on $\mathcal{P}$ can effectively guide AS remediation effort.
Figure 6: Comparison of Inspection Effort Analysis based on SAW Severity, $\mathcal{P}$, and AS Severity.

### v2

@@ 3.2 Study Context
We selected projects from the Qualitas Corpus (QCD) collection of software systems [91], using the compiled version of the QCD [90]. Table 3 details metrics on the QCD.
@@
Lines of Code (LOC) >> 18 million
@@ 3.3 Study setup and data collection
    We analyzed the QCD with three SATs and averaged the results for project packages;
@@
    To get our study results, we analyzed the finalized dataset, comprising all warning detected by the three SATs and the smells detected by the Arcan tool grouped by software package (see Section 3.4).
We analyzed the entire data set (RepoCount Java projects) from the Qualitas Corpus [91].
@@ 3.4 Data Analysis
We considered the packages affected by at least a smell and one static analysis warning at the class or package levels. One or more static analysis warnings could affect the same set of packages for each smell. This is because an smells may involve more than one class, while static analysis warnings involve only one class. In the case of projects not infected by static analysis warnings or smells, we did not consider them for the analysis.
To answer RQ1, we collected the QCD projects and ran the SAT and ARCAN. We merged the outputs of the tool. In our context, the dependent variable is the smell, and the independent variables are the warning. To compute the correlation coefficients, we need to assess the distributions of our variable. Therefore we conjectured one hypothesis ($H_{\mathcal{N}}$) as follows:
  • $H_{1\mathcal{N}}$
    The collected warning and smells metrics do not follow a normal distribution.
@@
  • $H_{0\mathcal{N}}$
    The collected warning and smells metrics follow a normal distribution.
We tested for normality ($H_{\mathcal{N}}$) using the Anderson–Darling (AD) test [86], considered one of the most powerful methods for detecting deviations from normality [87]. Since normality could not be assumed, we used Spearman’s $\rho$ [85] instead of Pearson’s correlation coefficient [70].
Table 4 presents $\rho$ values interpretation according to Dancey and Reidy [15].
@@
Therefore, we tested $H_{01(w,s)}$ with Spearman’s $\rho$ for each warning-smells pair, i.e., warning $w$ and smell $s$, resulting in 663 distinct tests. Each test measured the strength of correlation between the number of instances of $w$ and $s$ at the package level.
To answer RQ2, we computed Spearman’s $\rho$ for each warning-smells combination. We used the interquartile ratio (IQR) method [37] to select the top correlated warning for each SAT and smells, i.e., we computed the quartiles. We selected only the warnings with their Spearman’s $\rho$ value higher than or equal to the third quartile (Q3), thus selecting the top 25%. We computed the $\mathcal{P}$ as follows: Let $\Sigma_{\text{SAT}_{i}}$ represent the set of Static Analysis Warnings specific to a particular Static Analysis Tool $\text{SAT}_{i}$. Then, $\sigma\in\Sigma_{\text{SAT}_{i}}$ signifies that $\sigma$ belongs to the set of warnings specific to $\text{SAT}_{i}$; thus $\sigma_{j}$ is a specific warning and $\alpha\in smells$ represents a specific smells.. Therefore, we define the relative frequency $\mathcal{P}(\sigma_{j},\alpha)$ smells:
@@
  • The numerator represents the number of times the specific warning, $\sigma_{j}\in\Sigma_{\text{SAT}_{i}}$, appears within the given smells, $\alpha$.
  • The denominator represents the total count of all warnings belonging to a specific SAT, $\sigma$, in the given smells, $\alpha$.
Informally, $P(\sigma_{j},\alpha)$ expresses how likely a specific warning $\sigma_{j}$ is to appear within the context of packages affected by a given smell type $\alpha$. This helps identify which warnings are most frequently associated with particular architectural issues.
More specifically, we computed $\mathcal{P}$ for each possible combination of smells. Therefore, our data computed a $\mathcal{P}$-value for the 3 smells, the non-co-occurring instances, and for combinations of smells. We tested $H_{2}$, and $H_{3}$ via the WPT [14], which is a non-parametric statistical test that compares two related samples or paired data. WPT uses the absolute difference between the two observations to classify and then compare the sum of the positive and negative differences. The test statistic is the lowest of both. We selected WPT to test $H_{2}$ and $H_{3}$ because the data was not normally distributed; hence, we used it instead of the paired t-test, which assumes a normal data distribution.
Moreover, due to the high number of statistical tests performed, we needed to apply p-value correction. In place of the usual level of significance ($\alpha=0.05$), we need to correct p-values since it is essential in multiple hypothesis testing to control the increased risk of Type I errors. When multiple tests are conducted simultaneously, the probability of incorrectly rejecting at least one true null hypothesis increases. A classical p-value correction technique is the Bonferroni correction. However, Bonferroni is known to offer a conservative correction [36]. Therefore, we decided to use the Benjamini & Hochberg (BH) p-value correction [10], which uses the false discovery rate (FDR) when performing multiple hypothesis tests, smells follows [19]:
@@
    Rank the p-values: Sort the p-values obtained from multiple tests in ascending order, denoted smells $p_{(1)}\leq p_{(2)}\ldots\leq p_{(m)}$ for their respectively defined hypotheses $H_{(1)},H_{(2)},\ldots,H_{(m)}$.
@@
    where $m$ is the total number of tests, $q_{(i)}$ is the corrected probability for the hypothesis at moment $i$ and $p_{(j)}$ given the order rank $j\in\{1,\ldots,m\}$ of p-values, which stands smells the probability at the rank position $j$.
@@
    Determine correction index: The detailed procedure identified the most extensive index $k$ for which BH holds and rejects null hypotheses as follows:
@@
    Which is automatically computed using existing analytical software.
@@
    To determine which hypotheses to reject, to the desired critical value, $\alpha$ (e.g., 0.05).
Since we performed the BH correction on the p-values of the defined hypotheses, we maintained the standard value of the critical value $\alpha$ at 0.05, hence gaining power in controlling for error Type I and Type II [19].
Nevertheless, while BH correction mitigates false discoveries in large-scale testing, it may also suppress weaker, yet practically relevant, correlations. Future studies may benefit from exploring complementary methods such as smells, Bayesian correlation inference, or effect-size filtering to balance statistical rigor with exploratory sensitivity better.
To answer RQ3, we leverage our definition of $\mathcal{P}$, the default severity associated with the specific warning, and the severity associated with the particular smells to build three different rankers. The ranking idea stemmed from previous research by Çarka et al. [12], in which PofB and PopT metrics were normalized and used to evaluate classifier prediction concerning inspection effort (see Section 2.3).
@@
  • warning severity: we used SonarQube severity associated with each warning.
  • $\mathcal{P}$: we used $\mathcal{P}$ and created separate rankings one for each smells combination.
  • smells severity: we used the known smell severity to order the data.
We note that the smells severity-based order is a special case; by design, that ranker is the optimal one because it uses the smells severity. RQ3 aims to use the other two rankers to prioritize warnings that are more prone to induce smells. Therefore, the optima ranker is needed to compare the results of the other two. According to the interpretation of PofB and PopT, we selected the first x% of the data from the three rankings, with x ranging between 10 and 100 with increments of 10, e.g., top 10% ordered according to $\mathcal{P}$.
Therefore, for each selected percentage of data, we computed how many Critical, High, and Medium smell-prone warnings we would capture.
@@ 4.1 Static Analysis Warnings and Architectural Smell Correlation (RQ1)
We tested $H_{11}(w,s)$, excluding all cross-correlations among warnings and focused on the correlations between warning and smell (Figure 2), resulting in the rejection of the null hypothesis for each of the 663 warning-smell pair we measured the correlation of. Hence, we can affirm that there is a statistically significant correlation between the presence of static analyzer warnings and the presence of architectural smells. More specifically, warnings were weakly to moderately correlated with smell for each smell across the SAT.
Moreover, According to Table 5, we observe that the strongest correlations are associated with CheckStyle and FindBugs, with CheckStyle’s javadoc and sizes rules correlating with CD at $\rho$ values above 0.47, and FindBugs’ The Bloaters reaching a peak correlation of $\rho$ = 0.51 with CD. CD generally shows the highest correlations across tools, followed by HL and UD. PMD also shows consistent correlations, particularly for CD, with warnings such as Bean Members Should Serialize and Error Prone. These findings reinforce the idea that specific warnings, particularly those associated with structural or stylistic concerns, are statistically aligned with the presence of architectural erosion symptoms. Indeed, while the overall correlation is moderate, distinct and tool-specific warnings exhibit a stronger association with architectural smells.
Figure 2: Spearman’s $\rho$ average distribution for warning by smell and SAT (RQ1)
Table 5: Top 5 Most Significant Correlations Between warnings and smell by SAT
SAT WARNING SMELL $\rho$ SAT WARNING SMELL $\rho$
SonarQube squid:S1066 HL 0,250915 CheckStyle javadoc CD 0,485681
SonarQube squid:S1118 HL 0,248187 CheckStyle sizes CD 0,471169
SonarQube squid:S1155 HL 0,234255 CheckStyle Line Length Check CD 0,466939
SonarQube squid:S1226 HL 0,228857 CheckStyle checks CD 0,461382
SonarQube squid:S1135 HL 0,228094 CheckStyle Final Parameters Check CD 0,450387
SonarQube squid:Method Cyclomatic Complexity CD 0,223376 CheckStyle CD HL 0,574997
SonarQube squid:S134 CD 0,219365 CheckStyle javadoc HL 0,332039
SonarQube squid:S1066 CD 0,187805 CheckStyle checks HL 0,308395
SonarQube squid:S1188 CD 0,173311 CheckStyle The Bloaters HL 0,305625
SonarQube squid:S1226 CD 0,165946 CheckStyle Final Parameters Check HL 0,298404
SonarQube squid:Class Cyclomatic Complexity UD 0,191145 CheckStyle Overload Methods Declaration Order Check UD 0,212942
SonarQube squid:S1226 UD 0,141715 CheckStyle javadoc UD 0,207462
SonarQube squid:S1067 UD 0,139032 CheckStyle checks UD 0,206098
SonarQube squid:S134 UD 0,136966 CheckStyle The Bloaters UD 0,20556
SonarQube squid:S1066 UD 0,136421 CheckStyle Hide Utility Class Constructor Check UD 0,204744
SAT WARNING SMELL $\rho$ SAT WARNING SMELL $\rho$
FindBugs The Bloaters CD 0,510642 PMD Bean Members Should Serialize CD 0,488396
FindBugs The Dispensables CD 0,400668 PMD Error Prone CD 0,486998
FindBugs STYLE CD 0,341812 PMD Only One Return CD 0,482231
FindBugs PERFORMANCE CD 0,292369 PMD Code Style CD 0,471469
FindBugs The Object - Orientation Abusers CD 0,282612 PMD Documentation CD 0,465746
FindBugs The Bloaters HL 0,396364 PMD Documentation HL 0,319861
FindBugs The Dispensables HL 0,303299 PMD The Bloaters HL 0,309873
FindBugs The Encapsulators HL 0,250777 PMD Long Variable HL 0,305122
FindBugs code_smells:baseclass _abstract HL 0,228306 PMD Method Argument Could BeFinal HL 0,298117
FindBugs STYLE HL 0,218801 PMD Code Style HL 0,294548
FindBugs The Bloaters UD 0,257588 PMD Excessive Public Count UD 0,265235
FindBugs The Object - Orientation Abusers UD 0,254754 PMD GodClass UD 0,237967
FindBugs code_smells:baseclass _abstract UD 0,22347 PMD Excessive ClassLength UD 0,228839
FindBugs The Dispensables UD 0,211089 PMD Too Many Methods UD 0,221366
FindBugs The Encapsulators UD 0,206185 PMD Null Assignment UD 0,217341
@@ 4.2 Static Analysis Warnings inducing Architectural Smell (RQ2)
We tested $H_{2}$ with WPT. We rejected the null hypothesis for 243,258 out of 275,380 warning-smell pairs, i.e., 88%. Therefore, on average, we can affirm that there is a statistically significant difference in the co-occurrence of pairs of warning and smell. More specifically, Figure 3 presents the distribution of $H_{2}$ test results. It is worth noticing that in the case of UD, there was almost an equal probability of rejecting or accepting the null hypothesis across warnings. Moreover, we note that there is an equal probability of rejecting the null hypothesis in the non-co-occurring cases and those affected by all three smells.
Furthermore, we tested $H_{3}$ with WPT. We can reject the null hypothesis in 30 out of 42 cases, i.e., 72% of the cases. Table 6 shows the pair of SAT by smell presence for which we could not reject the null hypothesis. Therefore, we can affirm that there is a statistically significant difference in the co-occurrence of SAT-specific warnings and smell presence.
@@
Table 7 presents the top five warning co-occurring with specific smell. We selected the top five warnings that co-occur the most for each smell, including the “non-co-occurring”. We note that the most co-occurring warnings belong to FindBugs. Similarly, FindBugs also has the most “non-co-occurring” co-occurring warnings. Moreover, where three smells affect a single package, the FindBugs warning co-occurs each time. Furthermore, the least co-occurring smell is UD.
Table 7: The top five warning co-occurring with specific smell
SAT WARNING SMELL % co-occurring
@@
FindBugs NM_ METHOD_ NAMING_ CONVENTION NCO 89,865
@@ 4.3 Architectural Smell Prioritization (RQ3)
Figure 4 presents the distribution of $\mathcal{P}$ averaged across warnings across SAT per smell, i.e., the percentage of occurrences of warnings related to the occurrence of one or more smells or their absence (i.e., non-co-occurring). On average, warnings belonging to SonarQube are more likely to raise warnings about non-co-occurring packages and classes. Most warnings co-occur with CD smell and have a similar presence among the SATs. Moreover, second to CD, warnings co-occurred with all three smells. In this context, FindBugs’ warnings exhibit an average higher co-occurrence.
Figure 4: Distribution of $\mathcal{P}$ averaged among warning across SAT per smell (RQ2)
Figure 5 presents the warning prioritization based on warning Severity-based, $\mathcal{P}$-based, and optimal. We computed how many medium, high, or critical smell could be discovered by inspecting the first x% of data ranked by one of the three methods. According to its design, the optimal ranking prioritized the warning co-occurring with high and critical smell. The warning severity-based and UD+CD rankers demonstrate the most balanced performance, while the ranking based on warning co-occurring with CD favors moderately severe smell.
Finally, we tested $H_{4}$ with the WPT. We could reject the null hypothesis in all cases for high and critical smell severities, in which the optimal ranker was compared with the warning-severity-based and the $\mathcal{P}$ based rankers. Conversely, we could reject the hypothesis when comparing the warning-severity-based and the $\mathcal{P}$ based rankers, but only in the specific case when comparing the UD + HL + CD proneness based on $\mathcal{P}$. Furthermore, when the difference was statistically significant, the gains were in favor of the optimal rankers, as expected. Due to space constraints, we included the whole table and the graphical representation of the gains among the approaches in the replication package. Moreover, we can not reject the null hypothesis for the medium risk for each ranker. Therefore, the optimal ranker behaves similarly to the other two rankers.
@@
Nevertheless, based on the results, we can affirm that warning prioritization based on $\mathcal{P}$ can effectively guide smell remediation effort.
Figure 5: Comparison of warning prioritization based on warning Severity-based, $\mathcal{P}$-based, and optimal.